# Reinforcement Learning vs. Distillation: Understanding Accuracy and Capability in LLM Reasoning

### v1

@@ 5.2 Qualitative Analysis
In addtion, given that RLVR improves the model’s ability to generate quality responses, we examine how RLVR changes the qualitative characteristics of responses. Prior work suggests that RLVR-trained models often produce longer answers and show reflective reasoning behaviors such as verification and backtracking DeepSeek-AI (2025); Gandhi et al. (2025); Liu et al. (2025); Yeo et al. (2025); Zeng et al. (2025). Following these, we compare responses from the base and RL models along two surface-level dimensions: response length and the frequency of reflection-related keywords (e.g., “let’s verify,” “alternatively,” “wait”).
@@
6 Why Does Teacher Distillation Typically Increase Capability when RLVR Keeps It Unchanged?
@@ 6.1 Capability Analysis
Yue et al. briefly explored this issue by comparing two models: Qwen-2.5-Math-7B Yang et al. (2024) and DeepSeek-R1-Distill-Qwen-7B DeepSeek-AI (2025). The latter is a publicly available model obtained by distilling 800K DeepSeek-R1 responses into the Qwen-2.5-Math-7B student model. Their experiments showed that the distilled model demonstrates improved capability, as evidenced by higher pass@$k$ scores. However, it remains unclear why the distillation process led to the observed capability improvement. Teacher model responses contain two key components: (1) the model’s reasoning patterns, and (2) its domain knowledge. In the case of DeepSeek-R1-Distill-Qwen-7B, the student model was trained on up to 800K responses generated by DeepSeek-R1 DeepSeek-AI (2025), making it highly likely that new mathematical knowledge absent in the original student model’s pre-training data was also introduced during distillation. As a result, it is difficult to determine whether the observed gains in capability stem from adopting more effective reasoning pattern or from learning new knowledge.
@@
We train the reasoning-only model using the following procedure. Starting with the base model, we generate 256 responses for each of the 7,500 questions in the MATH train set. We exclude any question for which the model achieves a zero success rate. Since the base model answers each remaining question correctly at least once, we treat these as in-distribution. As the teacher, we use QwQ-32B Qwen (2024), which we verified to have higher capability than the student model (see Appendix A.7 for details). For each in-distribution question, QwQ-32B generates 8 candidate responses. From these, we randomly select up to 4 correct ones (using all available if fewer than 4 exist) and conduct SFT on the base model with them. Distilling in this manner should not introduce any new knowledge beyond what the base model already has. The resulting distilled model achieves 68% accuracy on the MATH 500 dataset, outperforming the base model’s 60%, indicating a successful distillation.
We evaluate the capability of the three models by conducting pass@$k$ experiments on the AIME25 OpenCompass (2025) and MATH500 dataset. Results are shown in Figure 4. We observe the following patterns consistently across both datasets. (1) The DeepSeek model outperforms the base model across the entire range of $k$ values. For AIME25, it achieves a pass@256 score of 70.0%, compared to 56.7% for the base model. To ensure this gap does not close at higher $k$, we extend evaluation of the base model up to pass@512 and confirm that its performance plateaus, remaining at 56.7%. A similar pattern is observed in MATH500. This shows that the DeepSeek model achieves gains in both accuracy and capability, consistent with (Yue et al., 2025). (2) In contrast, the reasoning-only model initially outperforms the base model at lower $k$, but the curves converge and even cross as $k$ increases. We also replicate the experiment with Qwen2.5-3B and observe the same convergence pattern (See Appendix A.8 for results). These results show that the reasoning-only model improves accuracy but does not expand capability, similar to the case of RLVR discussed in Section 4.
These results lead to following insight: Teacher distillation improves capability only when it introduces new knowledge. Learning reasoning patterns alone boosts accuracy but not capability.
@@ 6.2 In-Depth Analysis Based on Question Difficulty
These results suggest the following insight: Distillation with new knowledge improves both accuracy and capability because it enables the model to solve questions across all difficulty levels, including the most challenging ones. In contrast, reasoning-only distillation improves accuracy but not capability because, like RLVR, it focuses on easier questions—often at the cost of performance on the hardest ones.
@@ 7 Conclusions
Recent prior work has shown that RLVR improves accuracy but not capability, whereas teacher distillation from a strong teacher typically improves both accuracy and capability. In this paper, we provide explanations for these two phenomena, which are backed up with extensive empirical experimentation. In summary, the RLVR training algorithm focuses on improving the accuracy of the easier problems while paying less attention to the difficult problems; thus the difficult problems on the whole continue to have very low success rates, and thus the model’s overall capability remains unchanged. Teacher distillation is typically done with a strong teacher model which includes knowledge not present in the original base model; this leads to an increase in capability. Comparing RLVR and teacher distillation is thus an apples-to-oranges comparison, in that RLVR only uses the resources it has on hand, whereas teacher distillation augments the model with new resources.
@@
    We explain why RLVR improves accuracy but not capability, showing that it disproportionately favors easier questions to the detriment of the harder ones—often degrading performance for difficult questions.
@@
    We demonstrate that RLVR leads to better quality responses, as evidenced by self-distillation experiments, even though surface-level indicators such as response length or reflection keywords fail to capture this improvement.
@@
    We show that teacher distillation improves capability only when it introduces new knowledge, whereas distilling only reasoning patterns behaves similarly to RLVR—boosting accuracy at the expense of capability.
Together, these findings offer a clearer understanding of how RLVR and distillation shape reasoning behavior in LLMs.
@@ 8 Limitations
While our study offers an in-depth analysis of RLVR and distillation, several limitations point to future directions. First, our experiments are limited to relatively small models (1.5B and 3B), a single domain (mathematics), and a single RL algorithm family (GRPO & Dr. GRPO). Larger models, alternative domains, or different RL algorithms may exhibit different dynamics. A more comprehensive study across model scales, tasks, and training methods is needed to test the generality of our findings. Second, while we observe the importance of response quality in distillation across different experiments, the underlying factors that determine such quality remain underexplored. Future research may investigate what makes a response effective for improving accuracy and generalization, and propose more principled metrics for evaluating reasoning quality. Third, our results suggest that capability gains in distillation primarily stem from the introduction of new knowledge absent from the base model. While our approach offers a practical way to isolate this effect, future work could develop more rigorous frameworks to better characterize, quantify, and control knowledge transfer during distillation.
@@ A.8 Teacher Distillation Pass@$k$ results
In this paper, we used two models as base models: Qwen2.5-1.5B-Math and Qwen2.5-3B. For the RLVR-trained version of the 1.5B model, we used Qwen2.5-Math-1.5B-Oat-Zero [^2] https://huggingface.co/sail/Qwen2.5-Math-1.5B-Oat-Zero, a publicly available model trained by Liu et al.. According to their report, the model was trained with Dr.GRPO Liu et al. (2025), a variant of the GRPO algorithm Shao et al. (2024) designed to remove response length and question difficulty biases. The model was trained on questions from level 3 to 5 from the MATH training set. For the 3B model, we performed RLVR training ourselves. Training details are shown right below in Appendix A.10
@@ A.10.1 Prompt Setting
Prior work has shown that the performance of smaller models can be sensitive to prompt design Hochlehnert et al. (2025); Liu et al. (2025). Following Liu et al., we evaluated three prompt formats, as listed below. We ultimately adopted Template 3 (question only), which yielded the best performance.
@@ A.10.2 Reward Function
$R(q,a,r)=\begin{cases}1&\text{if the response $r$ to question $q$ matches the ground truth answer $a$}\\ -1&\text{otherwise}\end{cases}$

### v2

@@ 5.2 Qualitative Analysis
In addition, given that RLVR improves the model’s ability to generate quality responses, we examine how RLVR changes the qualitative characteristics of responses. Prior work suggests that RLVR-trained models often produce longer answers and show reflective reasoning behaviors such as verification and backtracking DeepSeek-AI (2025); Gandhi et al. (2025); Liu et al. (2025b); Yeo et al. (2025); Zeng et al. (2025). Following these, we compare responses from the base and RL models along two surface-level dimensions: response length and the frequency of reflection-related keywords (e.g., “let’s verify,” “alternatively,” “wait”).
@@
6 Under What Conditions Does Distillation Increase Capability?
@@ 6.1 Capability Analysis
Yue et al. briefly explored this issue by comparing two models: Qwen-2.5-Math-7B Yang et al. (2024) and DeepSeek-R1-Distill-Qwen-7B DeepSeek-AI (2025). The latter is a publicly available model obtained by distilling 800K DeepSeek-R1 responses into the Qwen-2.5-Math-7B student model. Their experiments showed that the distilled model demonstrates improved capability, as evidenced by higher pass@$k$ scores.
However, it remains unclear why the distillation process led to the observed capability improvement. Teacher model responses contain two key components: (1) the model’s reasoning patterns, and (2) its domain knowledge. In the case of DeepSeek-R1-Distill-Qwen-7B, the student model was trained on up to 800K responses generated by DeepSeek-R1 DeepSeek-AI (2025), making it highly likely that new mathematical knowledge absent in the original student model’s pre-training data was also introduced during distillation. As a result, it is difficult to determine whether the observed gains in capability stem from adopting more effective reasoning pattern or from learning new knowledge.
@@
We train the reasoning-only model using the following procedure. Starting with the base model, we generate 256 responses for each of the 7,500 questions in the MATH train set. We exclude any question for which the model achieves a zero success rate. Since the base model answers each remaining question correctly at least once, we treat these as in-distribution. As the teacher, we use QwQ-32B Qwen (2024), which we verified to have higher capability than the student model (see Appendix A.7 for details). For each in-distribution question, QwQ-32B generates 8 candidate responses. From these, we randomly select up to 4 correct ones (using all available if fewer than 4 exist) and conduct SFT on the base model with them. Distilling in this manner should not introduce any new knowledge beyond what the base model already has. The resulting distilled model achieves 69% accuracy on the MATH 500 dataset, significantly outperforming the base model’s 60%, indicating a successful distillation.
We evaluate the capability of the three models by conducting pass@$k$ experiments on the AIME25 OpenCompass (2025) and MATH500 dataset. Results are shown in Figure 4. We observe the following patterns consistently across both datasets. (1) The DeepSeek model outperforms the base model across the entire range of $k$ values. For AIME25, it achieves a pass@256 score of 70.0%, compared to 56.7% for the base model. To ensure this gap does not close at higher $k$, we extend evaluation of the base model up to pass@512 and confirm that its performance plateaus, remaining at 56.7%. A similar pattern is observed in MATH500. This shows that the DeepSeek model achieves gains in both accuracy and capability, consistent with (Yue et al., 2025). (2) In contrast, the reasoning-only model initially outperforms the base model at lower $k$, but the curves converge and even cross as $k$ increases. We also replicate the experiment with Qwen2.5-3B and observe the same convergence pattern (See Appendix A.8 for results). These results show that the reasoning-only model we train improves accuracy but does not expand capability, similar to the case of RLVR discussed in Section 4.
These results show that distillation does not always lead to capability expansion, even when it yields significant accuracy gains. We conjecture that the observed difference arises from whether new knowledge is introduced during distillation. Specifically, Distillation may improve capability when it introduces new knowledge, whereas learning reasoning patterns alone boosts accuracy but not capability.
@@ 6.2 In-Depth Analysis Based on Question Difficulty
We again conjecture that the key factor underlying this difference is whether new knowledge is introduced during distillation. Specifically, distillation with new knowledge improves both accuracy and capability because it enables the model to solve questions across all difficulty levels, including the most difficult ones. In contrast, reasoning-only distillation improves accuracy but not capability because, like RLVR, it focuses on easier questions—often at the cost of performance on the hardest ones. We hope this result motivates further empirical study to validate this conjecture and clarify the role of new knowledge in capability expansion.
@@ 7 Conclusions
Recent prior work has shown that RLVR improves accuracy but often fails to improve capability, whereas distillation from a strong teacher typically improves both. In this paper, we provide explanations for these two phenomena, supported by extensive empirical experiments. In summary, the RLVR training algorithm focuses on improving the accuracy of easier problems while paying less attention to difficult ones; thus, difficult problems on the whole continue to have very low success rates, and the model’s overall capability remains largely unchanged. Distillation, on the other hand, is typically performed with a strong teacher model that may include knowledge not present in the base model, which can lead to capability expansion.
@@
    We explain why RLVR improves accuracy but not capability by showing that it disproportionately favors easier questions to the detriment of harder ones—often degrading performance for difficult questions.
@@
    We demonstrate that RLVR yields higher-quality responses, as evidenced by self-distillation experiments, even though surface-level indicators such as response length or reflection keywords fail to capture this improvement.
@@
    While distillation consistently improves accuracy, its effect on capability is less clear. We conjecture that capability improves only when new knowledge is introduced, whereas distilling reasoning patterns only improves accuracy but not capability, sacrificing performance on the most difficult questions, similar to RLVR.
Together, these findings provide a clearer and more nuanced understanding of how RLVR and distillation shape reasoning behavior in LLMs.
@@ 8 Limitations
While our study provides an in-depth analysis of RLVR and distillation, it also has limitations that suggest directions for future work.
First, due to resource constraints, our experiments are restricted to a single domain, mathematics, and different patterns may emerge in other tasks. There remains an ongoing debate about whether RLVR truly improves capability. As discussed in Section 2, some studies argue that RLVR does not enhance capability in general mathematical settings where both training and test sets contain heterogeneous problems with uncontrolled knowledge and difficulty. Others, however, show that RLVR can indeed expand capability when sufficient training compute is available and when training and test sets are carefully controlled in terms of problem type and difficulty. Therefore, follow-up work is needed to unify these perspectives and develop a more comprehensive understanding of the phenomenon.
Second, our experiments are limited to relatively small models (1.5B and 3B) and a single RL algorithm family (GRPO & Dr. GRPO). Larger models or different RL algorithms may exhibit different dynamics. A more comprehensive study across model scales and training methods is needed to test the generality of our findings.
Third, our distillation experiments are limited in both scale and control. The DeepSeek model used for comparison is distilled on approximately 800k teacher responses and trained from a different base model, whereas our reasoning-only distillation relies on roughly 30k responses from the same model. In addition, when we extend the setup to include teacher responses to out-of-distribution (OOD) questions, we do not observe measurable capability improvement, possibly due to the small number of OOD samples or limited coverage of new knowledge. Consequently, we cannot conclusively determine whether capability gains depend on the introduction of new knowledge. Future work should validate this conjecture under more controlled settings.
9 Acknowledgement
We gratefully acknowledge the support of the Center for AI and Robotics (CAIR) at New York University Abu Dhabi for this research.
@@ A.5 Self-Distillation Results
As discussed in Section 5.1, the 1.5B model shows that self-distillation leads to overfitting: while training accuracy increases significantly, test accuracy improves only slightly or even declines. In contrast, distilling RL responses into the base model yields the strongest generalization improvement, raising test accuracy from 62.6% to 74.2%—surpassing both the base and RLVR-trained models. Similar trends are observed for the Qwen2.5-3B model as well: distilling RL responses into the base model again leads to the highest test accuracy, outperforming all other configurations. This consistent pattern across model sizes reinforces the interpretation that RLVR produces responses of higher quality. Taken together, these results suggest that distillation performance itself may serve as a useful proxy for evaluating the quality of model responses—potentially offering a more meaningful signal than surface-level indicators such as response length or syntactic heuristics.
@@ A.6.1 Response Length
(a) Mean response length grouped by correctness.
(b) Mean response length stratified by difficulty and correctness.
Figure 11: Comparison of response lengths between base and RL models for Qwen-2.5-1.5B-Math. RLVR did not increase verbosity, and correct answers tended to be shorter.
(a) Mean response length grouped by correctness.
(b) Mean response length stratified by difficulty and correctness.
Figure 12: Comparison of response lengths between base and RL models for Qwen-2.5-3B. RLVR did not increase verbosity, and correct answers tended to be shorter.
@@ A.8 Teacher Distillation Pass@$k$ results
In this paper, we used two models as base models: Qwen2.5-1.5B-Math and Qwen2.5-3B. For the RLVR-trained version of the 1.5B model, we used Qwen2.5-Math-1.5B-Oat-Zero [^2] https://huggingface.co/sail/Qwen2.5-Math-1.5B-Oat-Zero, a publicly available model trained by Liu et al.. According to their report, the model was trained with Dr.GRPO Liu et al. (2025b), a variant of the GRPO algorithm Shao et al. (2024) designed to remove response length and question difficulty biases. The model was trained on questions from level 3 to 5 from the MATH training set. For the 3B model, we performed RLVR training ourselves. Training details are shown right below in Appendix A.10
@@ A.10.1 Prompt Setting
Prior work has shown that the performance of smaller models can be sensitive to prompt design Hochlehnert et al. (2025); Liu et al. (2025b). Following Liu et al., we evaluated three prompt formats, as listed below. We ultimately adopted Template 3 (question only), which yielded the best performance.
@@ A.10.2 Reward Function
[…]